OpenStudy (anonymous):

Herb walks half the distance to school at a rate of 4 feed per second, and then doubles his rate for the rest of the way so that he won't be late for school. Find his average rate in feet per second for the whole distance A. 5 B. 5 1/3 C. 5 1/2 D. 6 E. 15

OpenStudy (anonymous):

avg = (8N feet + 8N feet) / 3N minutes 16N / 3N = 16/3 fps so is it 5 1/3

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

No you are right...

OpenStudy (anonymous):

i am not entirely sure myself

OpenStudy (anonymous):

But I am..

OpenStudy (anonymous):

alright thanks for the help

OpenStudy (anonymous):

Remember: If you cover a distance at the speed of x and you cover the same distance with the different speed of y, Then the average speed in covering the whole distance is: \[Average Speed = \frac{2xy}{x + y}\]

OpenStudy (anonymous):

ok